10. Nickel(II) chloride crystallizes with water molecules in its crystalline lattice as the hydrated

salt NiCl2•xH2O. If 5.000 g of the hydrated salt were placed in an oven and the water molecules were driven from the lattice, then the resulting “anhydrous” salt was cooled and re-weighed to be 2.726 g, what must be the value of x in NiCl2•xH2O ?
Answer #1

Steps to find out formula of hydrate

1.Find out mass of the water that has left the compound.

Mass of water=Mass of hydrate-mass of anhydrate

Mass of hydrate=5 g

mass of anhydrate=2.726 g

Mass of water=5 g-2.726 g=2.274 g

2.Convert mass water to moles

Moles of water=Mass of water/molar mass of water

=2.274 g/18 g per mole=0.126 moles

3.calculate the moles of nickel chloride in the hydrate=2.726/129.5=0.021 moles of Nickel chloride

4.Calculate water to nickel chloride ratio=moles of water / moles of nickel chloride

=0.126 moles/0.021 moles=6

So formula of hydrate NiCl2.6H2O